OpenStudy (anonymous):

Fan and Medal !Help! The answer I got was way off :( (m-2) - 5 = 8 - 2(m-4)

OpenStudy (anonymous):

you want all the m terms on one side so (m-2)+2(m-4)=8-4

OpenStudy (anonymous):

8-5 sorry!

OpenStudy (anonymous):

then expand and simplify so m-2+ 2m-8=3

OpenStudy (anonymous):

keep simplifying 3m-10=3

OpenStudy (anonymous):

3m=13 m=13/3

OpenStudy (anonymous):

I dont think thats the right answer... my book says 23/3 :( but I couldnt figure out y

OpenStudy (anonymous):

you're right i made a sign error

OpenStudy (anonymous):

(m-2) - 5 = 8 - 2(m-4) (m-2)+2(m-4)=8+5 m-2+2m-8=13 3m-10=13 3m=23 m=23/3

OpenStudy (anonymous):

got it! makes sense!! thank you! :)

OpenStudy (anonymous):

no problem!